Evaluate through constructing table of values


Lim In x 1-cos t/sint


t→0

1
Expert's answer
2022-03-04T09:43:16-0500

Solution


For the given limit


"\lim_{t\rightarrow 0}\frac{1-cost}{sint}"


For the function "f(t)=\frac{1-cos(t)}{sin(t)}"


The table of values when "t\rightarrow 0" (t approaches zero) is






We can see from the table,


"\lim_{t\rightarrow 0}\frac{1-cost}{sint}=0"